OpenStudy (anonymous):

the dimensions of a rectangular piece of cardboard are 12 inches by 8 inches. a tray is formed by cutting squares from the corners of the cardboard, folding up the sides, and taking the corners. write an equation V(w), for the volume of the tray as a function of the width of the squares, w, that are cut out of the corners. list the reasonable values for w.

OpenStudy (anonymous):

length of the tray = 12-2w width = 8 - 2w height = w so volume= w(12-2w)(8-2w)

OpenStudy (anonymous):

omg your awsome! if your good at geometry i have a few more geometry questions i need to post :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz close this question first by clickiing on good answer